show that the roots of the equation z^4-1=0 form a cyclic group of order 4.
1
Expert's answer
2020-04-24T17:48:43-0400

"Z^4-1=0\\
Z^4=1\\
Z^4=e^{2\pi n}\quad n=0,1,2,3\\
Z=e^{\frac{2\pi n}{4}}\quad n=0,1,2,3\\"

therefore,

"Z_1=1 \\
Z_2=e^{\frac{\pi}{2}}=j\\
Z_3=e^{\pi}=-1\\
Z_4=e^{\frac{3\pi}{2}}=-j\\"

"A=\{1,j,-1,-j\}=\{z^0,z^1,z^2,z^3\},\quad z=j.\\


\therefore A=<z>=\{z^k|k=0,1,2,3\}"


Therefore set A is a cyclic group of order 4.